What is the energy of a photon of yellow light with a frequency of 5.03 x 104 Hz?
Answer #1

here,

for yellow light

the frequency of photon ,f = 5.03 * 10^14 Hz

the energy of photon of yellow light , U = h * f

U = 6.634 * 10^-34 * 5.03 * 10^14 J

U = 3.34 * 10^-19 J

the energy of photon is yellow light is 3.34 * 10^-19 J